OpenStudy (anonymous):

darla is building a new desk. to make sure she had made a square corner she measures 6 ft from the corner along one edge and 4ft from the corner of the other edge. how long should the diagonal be between those two points if the corner is a right angle

OpenStudy (anonymous):

us the Pythagorean theorem \[a^{2}+b^{2} = c^{2}\] Where c = the diagonal

OpenStudy (anonymous):

\[c^2=6^2+4^2\]\[c=\sqrt{52}\]\[c=2\sqrt{13}\]

OpenStudy (anonymous):

wants this rounded as well how do you figure it out

OpenStudy (anonymous):

\[\sqrt{52}=\sqrt{2*2*13}=2\sqrt{13}\]if teacher want decimal you use calculator,if not just leave radical = 7.21110551 depend on teacher want round tenth or hundred

OpenStudy (anonymous):

round tenth = 7.2 round hundred = 7.21